# Hand1000: Generating Realistic Hands from Text with Only 1,000 Images

### v2

@@ Abstract
Text-to-image generation models have achieved remarkable advancements in recent years, aiming to produce realistic images from textual descriptions. However, these models often struggle with generating anatomically accurate representations of human hands. The resulting images frequently exhibit issues such as incorrect numbers of fingers, unnatural twisting or interlacing of fingers, or blurred and indistinct hands. These issues stem from the inherent complexity of hand structures and the difficulty in aligning textual descriptions with precise visual depictions of hands. To address these challenges, we propose a novel approach named Hand1000 that enables the generation of realistic hand images with target gesture using only 1,000 training samples. The training of Hand1000 is divided into three stages with the first stage aiming to enhance the model’s understanding of hand anatomy by using a pre-trained hand gesture recognition model to extract gesture representation. The second stage further optimizes text embedding by incorporating the extracted hand gesture representation, to improve alignment between the textual descriptions and the generated hand images. The third stage utilizes the optimized embedding to fine-tune the Stable Diffusion model to generate realistic hand images. In addition, we construct the first publicly available dataset specifically designed for text-to-hand image generation. Based on the existing hand gesture recognition dataset, we adopt advanced image captioning models and LLaMA3 to generate high-quality textual descriptions enriched with detailed gesture information. Extensive experiments demonstrate that Hand1000 significantly outperforms existing models in producing anatomically correct hand images while faithfully representing other details in the text, such as faces, clothing and colors.
@@ Introduction
In recent years, text-to-image generation models have achieved significant progress, such as Stable Diffusion (Rombach et al. 2022; Esser et al. 2024), Multimodal Large Language Model (MM-LLM) (Koh, Fried, and Salakhutdinov 2024), and DiT (Peebles and Xie 2023). However, these models often struggle when generating images involving human bodies, particularly hands. The generated images frequently exhibit distorted or misaligned body parts that do not conform to normal physical characteristics. As shown in Figure 1, this problem is especially pronounced with hands, which often appear blurry, with fingers twisted or interlaced, and an incorrect number of fingers. Despite capturing the color and texture of human hands, these generated images clearly display noticeable inaccuracies.
The underlying causes of this phenomenon can be attributed to two primary factors: (a) The inherent complexity of the human hand’s structure, along with the occlusions between fingers, means that the Stable Diffusion model lacks the necessary prior knowledge of the precise physical structure of human hands, resulting in its inability to generate anatomically correct hands. (b) When the text prompt includes information related to human hands (e.g., ”thumbs up” and ”phone call gesture”), the model struggles to associate this with corresponding hand images, indicating its inability to accurately comprehend text containing hand-related information and thus failing to produce realistic hand visual appearance.
@@
Inspired by text-based image editing works, we propose a novel method named Hand1000 to generate anatomically correct hands using only 1,000 images for each target hand gesture in three stages. To address the issue (a) from the previous paragraph, our proposed method enhances hand prior knowledge about correct hand anatomy by leveraging a pre-trained hand gesture recognition model (i.e., Mediapipe hands (Zhang et al. 2020)) to extract gesture representation in the first stage. To tackle issue (b), in the second stage, we optimize the text embedding for each image by integrating the hand gesture representation in the first stage to ensure alignment between text and hand image. This optimized embedding is then used to fine-tune the stable diffusion model to generate realistic hand images in the third stage. In addition, we construct a dataset specifically for text-to-hand-image generation by leveraging hand gesture recognition datasets (i.e. HaGRID dataset (Kapitanov et al. 2024)). Image captioning models(e.g. BLIP2 (Li et al. 2023b), PaliGemma (Beyer et al. 2024), and VitGpt2 (Mishra et al. 2024)) are first used to produce a textual description, and LLaMA3 model (Touvron et al. 2023a) is employed to enrich the textual description with hand gesture information. After training on 1000 images in our dataset, our Hand1000 significantly outperforms the Stable Diffusion model in generating correct hands for the same textual prompts, which is shown in Figure 1.
@@
(1) We propose a new method, Hand1000, that empowers the Stable Diffusion model with the ability to generate realistic hands using only 1,000 images. To the best of our knowledge, Hand1000 is the first text-to-image model capable of achieving accurate hand generation with such a small-scale dataset.
(2) By leveraging image captioning models and the LLaMA3, we construct the first publicly available dataset specifically designed for generating hand images from textual descriptions. This dataset aims to facilitate and advance research within the community on this task.
(3) We showcase the power of enhancing hand features and text optimization to generate anatomically correct and realistic hand images from text. Hand1000 achieves significant improvements compared with existing methods.
@@
Text-based image generation has long been a prominent research area, with classical methods including Generative Adversarial Networks (GANs) and their variants (Kang et al. 2023; Tao et al. 2022; Xu et al. 2018; Ye et al. 2021; Zhu et al. 2019; Wang et al. 2023, 2024; Li et al. 2023a; Ma et al. 2023), as well as stable diffusion models and their derivatives (Sauer et al. 2023; Saharia et al. 2022b; Dhariwal and Nichol 2021; Ho, Jain, and Abbeel 2020; Ho et al. 2022; Sohl-Dickstein et al. 2015; Yang et al. 2024; Ye et al. 2024; Yu et al. 2024). These methods leverage the capabilities of CLIP (Radford et al. 2021) and Transformer (Vaswani et al. 2017), empowering them to generate images from textual descriptions. In recent years, numerous related studies have emerged, achieving notable success in the field such as GigaGAN (Kang et al. 2023), Imagen (Saharia et al. 2022b), and SDXL Turbo (Sauer et al. 2023). With the advent of ChatGPT (Brown et al. 2020; Achiam et al. 2023) and the development of large language models(LLM) (Chowdhery et al. 2023; Touvron et al. 2023b; Team et al. 2023), several multimodal models for text-to-image generation have also emerged, such as DALL-E (Ramesh et al. 2021). Despite the notable advancements these methods have achieved in image generation, they frequently encounter significant issues when producing images that include human hands, often resulting in severe distortions or anomalies in hand representations.
@@
Despite the effectiveness of text-based diffusion models in generating various types of images, significant issues persist when it comes to rendering images containing human hands. Recent studies have begun to address this challenge. On one hand, some research focuses on resolving the issue through image editing techniques. For instance, HandRefiner (Lu et al. 2023) initially employs a deep learning model to identify hand regions within an image. It then crops these hand regions and utilizes ControlNet (Zhang, Rao, and Agrawala 2023) to attempt to produce a hand shape that adheres more closely to ergonomic standards. However, this approach has significant limitations. Firstly, it requires another text-to-image model to generate the initial image, as it cannot generate images independently based on text prompts. Secondly, if the initial image contains extremely distorted and chaotic hand representations, such as the second image in the first column of Figure 1, HandRefiner (Lu et al. 2023) is evidently unable to extract and correct the hand region. On the other hand, there are efforts aimed at directly producing accurate images of human hands. For example, HanDiffuser (Narasimhaswamy et al. 2024) reconstructs textual prompts into three-dimensional parameters and then embeds these parameters into subsequent processes to guide the generation of realistic hand images. This approach faces a significant challenge: the training process is exceedingly cumbersome. While text-to-image generation is typically a two-dimensional process, this method complicates it by introducing a 3D reconstruction phase to obtain precise parameters. Additionally, the above methods require an enormous training dataset, in the order of hundreds of thousands of samples, to achieve noticeable results. In contrast, our Hand1000 method only requires fine-tuning on a training set of 1,000 images to generate realistic hand images. Note that we cannot use Handiffuser (Narasimhaswamy et al. 2024) as a baseline due to the source code is not publicly available.
@@
Assume we have a set of images depicting a target hand gesture, such as ”phone call”. To extract hand features associated with this gesture, we feed the images into a gesture recognition model(i.e. Mediapipe hands (Zhang et al. 2020)) to obtain features from the final layer of the network. Subsequently, these features are averaged to obtain a Mean Hand Gesture Feature representation of the gesture, which is used for training in the following stages. Simultaneously, the Mean Hand Gesture Feature is also preserved and utilized during the inference phrase.
@@
In the second stage of training, the main objective is to integrate text embedding with hand gesture features to facilitate the diffusion model to generate realistic hand images. Given a textual description as input, CLIP text encoder is employed to obtain the Text Embedding. Subsequently, the text embedding is concatenated with the Mean Hand Gesture Feature obtained from the first stage. To fuse the text and hand gesture features, a fully connected (FC) layer is employed to map the concatenated embedding to the same dimension of text embedding, resulting in the Fused Embedding. Notably, the FC layer is kept frozen to reduce computational costs, as the embedding will undergo further optimization in Stage II. A linear fusion of the fused embedding with the original text embeddings is subsequently performed to obtain the Double Fused Embedding:
where is the hyperparameter, is Double Fused Embedding, is Text Embedding, and is Fused Embedding.
@@
In the final stage of training, the Optimized Embedding obtained in Stage II is used as a condition to fine-tune the Stable Diffusion model with the loss function specified in Equation 1. In line with image editing works (Kawar et al. 2023), it is essential to keep the Optimized Embedding frozen at this stage to preserve the hand gesture information integrated during the second phase. This freezing strategy ensures that the embedding remains consistently aligned with the hand features and continues to accurately match the hand images.
@@
This section details the specifics of the dataset created for training Hand1000, provides a qualitative and quantitative comparison between Hand1000 and the existing methods, and presents the ablation studies of our method.
@@
As no existing datasets specifically designed for text-to-hand image generation are publicly available, we construct a new dataset to bridge this gap. The process of the dataset curation is illustrated in Figure 4. Our dataset is built upon HaGRID (Kapitanov et al. 2024)), which is a comprehensive collection of various hand gestures. This dataset features clear hand images and includes a diversity of individuals, clothing, and colors, making it highly suitable for our training and evaluation purposes. We construct our datasets using six hand gestures—”phone call,” ”four,” ”like,” ”mute,” ”ok,” and ”palm” from HaGRID. 1,000 images are selected for each gesture for training and testing respectively. These hand gestures include ones with high finger separation (e.g., ”palm”), tightly closed fingers (e.g., ”mute”), and complex poses (e.g., ”ok”), providing a comprehensive assessment of the model’s capabilities.
@@
For each image in the training set, the first stage of training involves 10 epochs with a learning rate of 1e-3. The second stage consists of 20 epochs with a learning rate of 1e-6. Besides, We use Mediapipe hands (Zhang et al. 2020) as the hand gesture recognition model while the text encoder is the CLIP (Radford et al. 2021) text encoder. The sampling scheme is DDIM (Song, Meng, and Ermon 2020) and the optimizer is Adam (Kingma 2014).
@@
As shown in Table 2, our Hand1000 model demonstrates a significant performance improvement over the Stable Diffusion (Rombach et al. 2022) in terms of all the metrics. Specifically, Hand1000 achieves a 28.599 reduction in FID, a 0.017 reduction in KID, a 30.208 reduction in FID-H, and a 0.041 reduction in KID-H, while its HAND-CONF score is elevated by 0.053. To further validate the effectiveness of our method, we also perform comparison with fine-tuned stable diffusion using the same dataset. Although the fine-tuned Stable Diffusion model demonstrates slightly superior performance in terms of FID and KID compared to our method, several studies (Kynkäänniemi et al. 2019; Borji 2019) have suggested that these two metrics are not entirely reliable. Nevertheless, the quality of hand image generation is significantly inferior, as indicated by FID-H, KID-H, and HAND-CONF. Compared to the fine-tuned stable diffusion model, our method achieves a 44.294 reduction in FID-H, a 0.027 reduction in KID-H, and a 0.035 increase in HAND-CONF, underscoring the effectiveness of our approach.
@@
By providing identical text prompts, we conduct a comparative analysis of the results with Stable Diffusion, finetuned Stable Diffusion as well as using Imagic (Kawar et al. 2023) to edit the results of Stable Diffusion (Stable Diffusion+Imagic). As illustrated in Figure 5, our method not only generates realistic hand images but also effectively renders other components such as characters and clothing, demonstrating superior performance compared to all the other methods. Furthermore, though Imagic is adopted to edit the image generated by Stable Diffusion, the result led to subtle alterations, such as the lightening of the dress color in the first set of images in the bottom row. In other words, Imagic fails to recover realistic hands and bodies from the generated images, leaving them in a chaotic, deformed, and distorted state. Additional experimental results can be found in the supplementary materials.
@@
The ablation results for the feature fusion and embedding optimization in Stage II are shown in Table 3. Since FID-H and KID-H metrics are calculated after extracting the hand region, they provide a more accurate measure of the model’s hand generation capability. With the addition of embedding optimization, FID-H decreased by 25.003 and KID-H by 0.035. Incorporating Feature Fusion, which integrates features extracted by the hand gesture recognition model with text embeddings, led to a reduction of 20.319 in FID-H and 0.048 in KID-H. When both techniques are used together, FID-H decreased by 29.636 and KID-H by 0.051, demonstrating the superiority of our designs.
@@
To determine the optimal training set size for our method, we conduct ablation experiments using FID-H as the evaluation metric. As shown in Figure 7, the experiments reveal that when the training set size is below 1,000, the FID-H is relatively high. The FID-H begins to decrease at a training set size of 500, reaching its minimum value at 1,000. At a training set size of 1,200, the FID-H shows a slight increase, but the magnitude of this increase is small. Therefore, we select 1,000 as the optimal training set size, and ultimately decided to use only 1,000 images for all experiments to ensure both efficiency and performance.

### v3

@@ Abstract
Text-to-image generation models have achieved remarkable advancements in recent years, aiming to produce realistic images from textual descriptions. However, these models often struggle with generating anatomically accurate representations of human hands. The resulting images frequently exhibit issues such as incorrect numbers of fingers, unnatural twisting or interlacing of fingers, or blurred and indistinct hands. These issues stem from the inherent complexity of hand structures and the difficulty in aligning textual descriptions with precise visual depictions of hands. To address these challenges, we propose a novel approach named Hand1000 that enables the generation of realistic hand images with target gesture using only 1,000 training samples. The training of Hand1000 is divided into three stages with the first stage aiming to enhance the model’s understanding of hand anatomy by using a pre-trained hand gesture recognition model to extract gesture representation. The second stage further optimizes text embedding by incorporating the extracted hand gesture representation, to improve alignment between the textual descriptions and the generated hand images. The third stage utilizes the optimized embedding to fine-tune the Stable Diffusion model to generate realistic hand images. In addition, we construct the first publicly available dataset specifically designed for text-to-hand image generation. Based on the existing hand gesture recognition dataset, we adopt advanced image captioning models and LLaMA3 to generate high-quality textual descriptions enriched with detailed gesture information. Extensive experiments demonstrate that Hand1000 significantly outperforms existing models in producing anatomically correct hand images while faithfully representing other details in the text, such as faces, clothing and colors. Additional details and resources are available on our project page: https://haozhuo-zhang.github.io/Hand1000-project-page/.
@@ Introduction
In recent years, text-to-image generation models have achieved significant progress, such as Stable Diffusion (Rombach et al. 2022; Esser et al. 2024), Multimodal Large Language Model (MM-LLM) (Koh, Fried, and Salakhutdinov 2024), and DiT (Peebles and Xie 2023). However, these models often struggle when generating images involving human bodies, particularly hands. The generated images frequently exhibit distorted or misaligned body parts that do not conform to normal physical characteristics. As shown in the top of Figure 1, this problem is especially pronounced with hands, which often appear blurry, with fingers twisted or interlaced, and an incorrect number of fingers. Despite capturing the color and texture of human hands, these generated images clearly display noticeable inaccuracies.
The underlying causes of this phenomenon can be attributed to two primary factors: (a) The inherent complexity of the human hand’s structure, along with the occlusions between fingers, means that the Stable Diffusion model lacks the necessary prior knowledge of the precise physical structure of human hands, resulting in its inability to generate anatomically correct hands. (b) When the text prompt includes information related to human hands (e.g., “thumbs up” and “phone call gesture”), the model struggles to associate this with corresponding hand images, indicating its inability to accurately comprehend text containing hand-related information and thus failing to produce realistic hand visual appearance.
@@
Inspired by text-based image editing works, we propose a novel method named Hand1000 to generate anatomically correct hands using only 1,000 images for each target hand gesture in three stages. To address the issue (a) from the previous paragraph, our proposed method enhances hand prior knowledge about correct hand anatomy by leveraging a pre-trained hand gesture recognition model (i.e., Mediapipe hands (Zhang et al. 2020)) to extract gesture representation in the first stage. To tackle issue (b), in the second stage, we optimize the text embedding for each image by integrating the hand gesture representation in the first stage to ensure alignment between text and hand image. This optimized embedding is then used to fine-tune the stable diffusion model to generate realistic hand images in the third stage. In addition, we construct a dataset specifically for text-to-hand-image generation by leveraging hand gesture recognition datasets (i.e. HaGRID dataset (Kapitanov et al. 2024)). Image captioning models(e.g. BLIP2 (Li et al. 2023b), PaliGemma (Beyer et al. 2024), and VitGpt2 (Mishra et al. 2024)) are first used to produce a textual description, and LLaMA3 model (Touvron et al. 2023a) is employed to enrich the textual description with hand gesture information. After training on 1,000 images in our dataset, our Hand1000 significantly outperforms the Stable Diffusion model in generating correct hands for the same textual prompts, which is shown at the bottom of Figure 1.
@@
We propose a new method, Hand1000, that empowers the Stable Diffusion model with the ability to generate realistic hands using only 1,000 images. To the best of our knowledge, Hand1000 is the first text-to-image model capable of achieving accurate hand generation with such a small-scale dataset.
By leveraging image captioning models and the LLaMA3, we construct the first publicly available dataset specifically designed for generating hand images from textual descriptions. This dataset aims to facilitate and advance research within the community on this task.
We showcase the power of enhancing hand features and text optimization to generate anatomically correct and realistic hand images from text. Hand1000 achieves significant improvements compared with existing methods.
@@
Text-based image generation has long been a prominent research area, with classical methods including Generative Adversarial Networks (GANs) and their variants (Kang et al. 2023; Tao et al. 2022; Xu et al. 2018; Zhu and Ngo 2020; Ye et al. 2021; Zhu et al. 2019b, a; Wang et al. 2024; Li et al. 2023a; Ma et al. 2023), as well as stable diffusion models and their derivatives (Sauer et al. 2023; Saharia et al. 2022b; Dhariwal and Nichol 2021; Ho, Jain, and Abbeel 2020; Ho et al. 2022; Sohl-Dickstein et al. 2015; Yang et al. 2024; Ye et al. 2024; Yu et al. 2024). These methods leverage the capabilities of CLIP (Radford et al. 2021) and Transformer (Vaswani et al. 2017), empowering them to generate images from textual descriptions. In recent years, numerous related studies have emerged, achieving notable success in the field such as GigaGAN (Kang et al. 2023), Imagen (Saharia et al. 2022b), and SDXL Turbo (Sauer et al. 2023). With the advent of ChatGPT (Brown et al. 2020; Achiam et al. 2023) and the development of large language models(LLM) (Chowdhery et al. 2023; Touvron et al. 2023b; Team et al. 2023), several multimodal models for text-to-image generation have also emerged, such as DALL-E (Ramesh et al. 2021). Despite the notable advancements these methods have achieved in image generation, they frequently encounter significant issues when producing images that include human hands, often resulting in severe distortions or anomalies in hand representations.
@@
Text-based diffusion models are effective for generating various images but struggle with rendering human hands. Recent works address this challenge in two ways. Some studies use image inpainting, like HandRefiner (Lu et al. 2023), which identifies and crops hand regions, then refines them using ControlNet (Zhang, Rao, and Agrawala 2023). However, this approach depends on another model for initial image generation and fails with severely distorted hands. Alternatively, methods like HanDiffuser (Narasimhaswamy et al. 2024) directly generate hand images by reconstructing text prompts into 3D hand parameters, though the training process is complex and expensive. Additionally, the above methods require an enormous training dataset, in the order of hundreds of thousands of samples, to achieve noticeable results. In contrast, our Hand1000 method fine-tunes on just 1,000 images to generate realistic hands without requiring explicit hand pose modeling, offering a more efficient and flexible approach.
@@
Assume we have a set of images depicting a target hand gesture, such as “phone call”. To extract hand features associated with this gesture, we feed the images into a gesture recognition model(i.e. Mediapipe hands (Zhang et al. 2020)) to obtain features from the final layer of the network. Subsequently, these features are averaged to obtain a Mean Hand Gesture Feature representation of the gesture, which is used for training in the following stages. Simultaneously, the Mean Hand Gesture Feature is also preserved and utilized during the inference phase.
@@
In the second stage of training, the main objective is to integrate text embedding with hand gesture features to facilitate the diffusion model to generate realistic hand images. Given a textual description as input, CLIP text encoder is employed to obtain the Text Embedding. Subsequently, the text embedding is concatenated with the Mean Hand Gesture Feature obtained from the first stage. To fuse the text and hand gesture features, a fully connected (FC) layer is employed to map the concatenated embedding to the same dimension of text embedding, resulting in the Fused Embedding, which smoothly incorporates the gesture’s general characteristics. Notably, the FC layer is kept frozen to reduce computational costs, as the embedding will undergo further optimization in Stage II. A linear fusion of the fused embedding with the original text embeddings is subsequently performed to obtain the Double Fused Embedding:
where is the hyperparameter, is Double Fused Embedding, is Text Embedding, and is Fused Embedding. This linear fusion reinforces the text embedding’s unique features, allowing it to leverage the gesture feature without overpowering its own nuances.
@@
In the final stage of training, the Optimized Embedding obtained in Stage II is used as a condition to fine-tune the Stable Diffusion model with the loss function specified in Equation 1. The fine-tuning procedure follows the standard text-to-image diffusion model training, with the only modification being the replacement of the usual text embedding with our Optimized Embedding from Stage II. To preserve the hand gesture information integrated in Stage II, the Optimized Embedding is kept frozen during this stage, in line with image editing works (Kawar et al. 2023).
@@
As no existing datasets specifically designed for text-to-hand image generation are publicly available, we construct a new dataset to bridge this gap. The process of the dataset curation is illustrated in Figure 4. Our dataset is built upon HaGRID (Kapitanov et al. 2024)), which is a comprehensive collection of various hand gestures. This dataset features clear hand images and includes a diversity of individuals, clothing, and colors, making it highly suitable for our training and evaluation purposes. We construct our datasets using six hand gestures—“phone call,” “four,” “like,” “mute,” “ok,” and “palm” from HaGRID. 1,000 images are selected for each gesture for training and testing respectively. These hand gestures include ones with high finger separation (e.g., “palm”), tightly closed fingers (e.g., “mute”), and complex poses (e.g., “ok”), providing a comprehensive assessment of the model’s capabilities.
@@
For each image in the training set, the first stage of training involves 10 epochs with a learning rate of 1e-3. The second stage consists of 20 epochs with a learning rate of 1e-6. Besides, We use Mediapipe hands (Zhang et al. 2020) as the hand gesture recognition model while the text encoder is the CLIP (Radford et al. 2021) text encoder. The sampling scheme is DDIM (Song, Meng, and Ermon 2020) and the optimizer is Adam (Kingma 2014). We used SDv1.4 for our experiments, which is similar to recent works like HanDiffuser (Narasimhaswamy et al. 2024), HandRefiner (Lu et al. 2023), and Imagic (Kawar et al. 2023).
@@
As shown in Table 2, our Hand1000 model demonstrates a significant performance improvement over the Stable Diffusion (Rombach et al. 2022) in terms of all the metrics. Specifically, Hand1000 achieves a 28.599 reduction in FID, a 0.017 reduction in KID, a 30.208 reduction in FID-H, and a 0.041 reduction in KID-H, while its HAND-CONF score is elevated by 0.053. To further validate the effectiveness of our method, we also perform comparison with fine-tuned stable diffusion using the same dataset. Although the fine-tuned Stable Diffusion model demonstrates slightly superior performance in terms of FID and KID compared to our method, several studies (Kynkäänniemi et al. 2019; Borji 2019) have suggested that these two metrics are not entirely reliable. Nevertheless, the quality of hand image generation is significantly inferior, as indicated by FID-H, KID-H, and HAND-CONF. Compared to the fine-tuned stable diffusion model, our method achieves a 44.294 reduction in FID-H, a 0.027 reduction in KID-H, and a 0.035 increase in HAND-CONF, underscoring the effectiveness of our approach. Furthermore, we also conduct experiments on the open-source hand image refinement method, HandRefiner (Lu et al. 2023), and the results demonstrate that our method outperforms HandRefiner (Lu et al. 2023) across all metrics, with an 11.396 reduction in FID, a 0.006 reduction in KID, a 10.037 reduction in FID-H, a 0.003 reduction in KID-H, and a 0.029 increase in HAND-CONF.
@@
By providing identical text prompts, we conduct a comparative analysis of the results with Stable Diffusion, finetuned Stable Diffusion, using Imagic (Kawar et al. 2023) to edit the results of Stable Diffusion and fine-tuned Stable Diffusion as well as HandRefiner. As illustrated in Figure 5, our method not only generates realistic hand images but also effectively renders other components such as characters and clothing, demonstrating superior performance compared to all the other methods. Furthermore, though Imagic is adopted to edit the image generated by Stable Diffusion, the result led to subtle alterations, such as the lightening of the dress color in the first set of images in the bottom row. In other words, Imagic fails to recover realistic hands and bodies from the generated images, leaving them in a chaotic, deformed, and distorted state. Additional experimental results can be found in the supplementary materials. Besides, HandRefiner fails to interpret gesture information accurately, despite some improvement in hand deformation, and shows no effect when only partial hands are present, as in the rightmost example.
@@
The ablation results for the feature fusion and embedding optimization in Stage II are shown in Table 3. Since FID-H and KID-H metrics are calculated after extracting the hand region, they provide a more accurate measure of the model’s hand generation capability. With the addition of embedding optimization, FID-H is decreased by 25.003 and KID-H by 0.035. Incorporating Feature Fusion, which integrates features extracted by the hand gesture recognition model with text embeddings, leads to a reduction of 20.319 in FID-H and 0.048 in KID-H. When both techniques are used together, FID-H is decreased by 29.636 and KID-H by 0.051, demonstrating the superiority of our designs.
@@
To determine the optimal training set size for our method, we conduct ablation experiments using FID-H as the evaluation metric. As shown in Figure 7, the experiments reveal that when the training set size is below 1,000, the FID-H is relatively high. The FID-H begins to decrease at a training set size of 500, reaching its minimum value at 1,000. At a training set size of 1,200, the FID-H shows a slight increase, but the magnitude of this increase is small, which may indicate overfitting of the embedding, a known challenge in image-editing methods and worth further study. Therefore, we select 1,000 as the optimal training set size, and ultimately decided to use only 1,000 images for all experiments to ensure both efficiency and performance.